# LoSiA: Efficient High-Rank Fine-Tuning via Subnet Localization and Optimization

### v1

@@ 1 Introduction
To answer this question, we propose LoSiA (Low-Resources Subnet Integration Adaptation), a novel PEFT framework that dynamically localizes and optimizes critical sub-networks periodically, as illustrated in Figure 1. LoSiA asynchronously selects a core sub-network for each layer by calculating sensitivity-based importance scores and performing greedy selecting algorithms. Following localization, it fine-tunes the identified sub-network and applies a rewarming learning rate strategy to promote stable and robust training. The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, which reduces training latency and ensures no steep increase in training time for high-rank updating. Additionally, LoSiA does not introduce extra architectural components and only requires an optimizer replacement for seamless deployment. Extensive experiments demonstrate its superior performance among PEFT baselines on domain-specific, commonsense reasoning tasks, while mitigating forgetting in continue learning. We also propose LoSiA-Pro, a more refined equivalent implementation of LoSiA, which significantly reduces the activation storage and computational complexity in backward propagation. LoSiA-Pro speeds up training 1.38×1.38$\times$1.38 × compared to LoRA and 2.68×2.68$\times$2.68 × compared to DoRA.
@@
(3) We conduct extensive evaluations across multiple models and benchmarks. LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by 1.15×1.15$\times$1.15 × compared to LoRA. Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of 1.38×1.38$\times$1.38 ×.
@@ 3 Method
[…]
@@ 3.1 Structure of Gradients
To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer: The 32 attention heads exhibit a significant disparity in attention scores, while gradients strongly correlate with output neurons $Y_{S}$. In particular, a consistent subset of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads. The sparse pattern remains consistent in MLP layers (Appendix A.2.1). We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ termed the core subnet - rather than the entire network.
@@ 3.2 Subnet Localization
[…]
@@
[…]
@@ 3.3 Subnet Optimization and Intergration
Naive periodic re-localization strategies can induce training instability and loss spikes Lialin et al. (2023). Furthermore, the storage requirements for $\overline{I}(\cdot),\overline{U}(\cdot)$ in a synchronous period would lead to a scaling of GPU memory overhead. Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate. Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each layer $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$ with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$. The training timeline is divided into time slots of $T$ steps, such that for time slots $[iT,(i+1)T),i=1,2,\ldots$, we:
@@
    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $[(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
@@
$\overline{lr}(t)=\begin{cases}\frac{t-(kL+l)T}{T}\cdot\text{lr}(t)&\text{if C}\\ \text{lr}(t)&\text{otherwise}\\ \end{cases}\$ (8)
The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}$, where $T_{w}$ is the warmup duration. Figure 4 illustrates the timelines of importance calculation and the learning rate rewarming across layers, with re-localization sandwiched between them.
@@ 4.3 Ablation Study
Synchronous Localization (SL) $42.76$ $44.13$ $43.45$
Gradient-based Localization (GL) $43.00$ $44.88$ $43.94$
w/o Warm-up during Selection (WDS) $38.06$ $44.21$ $41.14$
w FFT lm_head (FFTO) $43.96$ $44.32$ $44.14$
@@ 4.4 Analysis
[…]
@@
Seq-LoRA $66.62$ $\mathbf{1.46}$ $-8.04$
Seq-LoSiA $\mathbf{70.48}$ $-0.20$ $\mathbf{-3.54}$
@@ A.3.1 Domain Specific Tasks
Scale Related $\alpha=128$ $\alpha=2.0$
@@
Implement Layer proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj, lm_head
@@ A.3.2 Common-Sense Reasoning Tasks
Scale Related $\alpha=128$ $\alpha=2.0$
@@
Implement Layer proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj, lm_head
@@ A.3.3 Rank Factor Robustness
Sensitivity-based Localization (LoSiA) $41.70$ $\mathbf{52.23}$ $\mathbf{50.89}$ $\mathbf{36.82}$ $\mathbf{44.95}$
Gradient-based Localization (GL) $\mathbf{42.64}$ $51.41$ $50.62$ $36.22$ $44.88$
@@ A.3.4 Continue Learning
To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks by the order HellaSwag, PIQA, BoolQ, SIQA and Winogrande. Learning rate for LoRA is $41e-$ 4 and for LoSiA is $55e-$ 5. The remaining hyperparameters are consistent with Table 9. LoRA modules are merged into the backbone before subsequent task adaptation.
@@
Method Task (#1) HellaS (#2) PIQA (#3) BoolQ (#4) SIQA (#5) WinoG ST
Seq-LoRA HellaSwag $59.86$ $55.64$ $59.10$ $57.86$ $54.36$ $59.86$
PIQA $76.01$ $80.52$ $77.86$ $78.73$ $77.64$ $79.33$
BoolQ $77.80$ $73.27$ $86.30$ $80.12$ $75.93$ $88.07$
SIQA $45.80$ $47.80$ $45.85$ $59.52$ $46.11$ $56.86$
Winogrande $64.25$ $68.35$ $68.82$ $69.93$ $79.08$ $73.88$
Seq-LoSiA HellaSwag $63.72$ $61.89$ $61.11$ $60.37$ $\mathbf{56.43}$ $63.72$
PIQA $78.29$ $79.49$ $79.82$ $79.38$ $\mathbf{77.75}$ $81.50$
BoolQ $77.52$ $70.76$ $83.24$ $82.54$ $\mathbf{81.99}$ $84.13$
SIQA $47.80$ $48.26$ $48.26$ $59.93$ $\mathbf{56.04}$ $61.05$
Winogrande $68.51$ $67.88$ $68.51$ $71.82$ $\mathbf{80.19}$ $77.19$
@@ A.4.1 Memory Estimate
#Total $8LKrdb$ $2(LKR^{2}b+Vdb)$ $+\max\{d^{2}b,Vdb\}$ $+2LKRdb$ $2(LKd^{2}p^{2}b+Vdp_{o}b)$ $+\max\{d^{2}b,Vdb\}$ $+2Kd^{2}b$
@@ A.4.2 Latency Measurement
[…]

### v3

@@ 1 Introduction
To answer this question, we propose LoSiA (Low-Resources Subnet Integration Adaptation), a novel PEFT framework that dynamically localizes and optimizes critical sub-networks periodically, as illustrated in Figure 1. LoSiA asynchronously selects a core sub-network for each layer by calculating sensitivity-based importance scores and performing greedy selecting algorithms. Following localization, it fine-tunes the identified sub-network and applies a rewarming learning rate strategy to promote stable and robust training. The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, which reduces training latency and ensures no steep increase in training time for high-rank updating. Additionally, LoSiA does not introduce extra architectural components and only requires an optimizer replacement for seamless deployment. Extensive experiments demonstrate its superior performance among PEFT baselines on domain-specific, commonsense reasoning tasks, while mitigating forgetting in continue learning. We also propose LoSiA-Pro, a more refined equivalent implementation of LoSiA, which significantly reduces the activation storage and computational complexity in backward propagation. LoSiA-Pro speeds up training $1.38\times$ compared to LoRA and $2.68\times$ compared to DoRA.
@@
(3) We conduct extensive evaluations across multiple models and benchmarks. LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by $1.15\times$ compared to LoRA. Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of $1.38\times$.
@@ 3 Method
Consider a model $f_{0}:\mathcal{X}\to\mathcal{Y}$ trained on dataset $\mathcal{D}=\{B_{i}\}_{i=1}^{N}$, where each batch $B_{i}=\{(x_{j},y_{j})\}_{j=1}^{M}$ contains $M$ samples. Let $W$ denote the parameters and $\mathcal{L}$ the loss function. The neural network $S_{0}$ in $f_{0}$ can be represented as a tuple comprising input neurons $X_{S_{0}}$, output neurons $Y_{S_{0}}$ and neural connections $W_{X_{S_{0}},Y_{S_{0}}}$, that is, $S_{0}=(X_{S_{0}},Y_{S_{0}},W_{X_{S_{0}},Y_{S_{0}}})$. The notation $f_{0}\xrightarrow[\mathcal{D}]{P_{0}}f$ denotes training model $f_{0}$ on $\mathcal{D}$ over full parameters $P_{0}=\{W_{X_{S_{0}},Y_{S_{0}}}\}$ and produces model $f$. We investigate the following question: Can we efficiently identify a parameter subset $P\subset P_{0}$ of fixed size, such that training $f\xrightarrow[\mathcal{D}]{P}f^{\prime}$ minimizes the loss difference $\Delta\mathcal{L}=|\mathcal{L}(f^{\prime},\mathcal{D})-\mathcal{L}(f,\mathcal{D})|$?
@@ 3.1 Structure of Gradients
To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer: The 32 attention heads exhibit a significant disparity in attention scores, while gradients strongly correlate with output neurons $Y_{S}$. In particular, a consistent subset of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads. The sparse pattern remains consistent in MLP layers (Appendix A.2.1). We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ - termed the core subnet - rather than the entire network.
@@ 3.2 Subnet Localization
3:  $\text{rows}\leftarrow\text{Top-K}(\text{sums},\lfloor np\rfloor).\text{indices}$
@@
5:  $\text{cols}\leftarrow\text{Top-K}(\text{sums},\lfloor mp\rfloor).\text{indices}$
6:  return $(\text{rows},\text{cols})$
@@ 3.3 Subnet Optimization and Intergration
Naive periodic re-localization strategies can induce training instability and loss spikes Lialin et al. (2023). Furthermore, the storage requirements for $\overline{I}(\cdot),\overline{U}(\cdot)$ in a synchronous period would lead to a scaling of GPU memory overhead. Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate. Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each layer $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$ with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$. The training timeline is divided into time slots of $T$ steps, such that for time slots $iT,(i+1)T),i=1,2,\ldots$, we:
@@
    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
@@
[…]
The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}\$, where $T_{w}$ is the warmup duration. Figure 4 illustrates the timelines of importance calculation and the learning rate rewarming across layers, with re-localization sandwiched between them.
@@ 4.3 Ablation Study
Synchronous
Localization (SL)
$42.76$ $44.13$ $43.45$
Gradient-based
Localization (GL)
$43.00$ $44.88$ $43.94$
w/o Warm-up
during Selection (WDS)
$38.06$ $44.21$ $41.14$
w FFT lm_head
$43.96$ $44.32$ $44.14$
@@ 4.4 Analysis
Low-rank fine-tuning methods often introduce intruder dimensions Shuttleworth et al. (2024b), resulting in spectral discrepancies between the fine-tuned and the pre-trained weights. This diminishes the adaptability of LoRA in sequential learning. Figure 8 illustrates the cosine similarity between the $\text{Top-}500$ singular vectors of the trained matrices and those of the original weights. Both LoRA and DoRA exhibit dimensional shifts due to their low-rank structures, whereas LoSiA demonstrates higher similarity and dimensional stability comparable to FFT.
@@
Seq-LoRA $66.62$ $\mathbf{1.46}$ 8.04-8.04
Seq-LoSiA $\mathbf{70.48}$ 0.20-0.20 $\mathbf{-3.54}$
@@ A.3.1 Domain Specific Tasks
Scale Related $\alpha=128$ $\alpha=2.0$ -
@@
Implement Layer
proj_q,proj_k,proj_v,proj_o,
up_proj,down_proj,gate_proj
proj_q,proj_k,proj_v,proj_o,
up_proj,down_proj,gate_proj,
lm_head
@@ A.3.2 Common-Sense Reasoning Tasks
Scale Related $\alpha=128$ $\alpha=2.0$ -
@@
Implement Layer
proj_q,proj_k,proj_v,proj_o,
up_proj,down_proj,gate_proj
proj_q,proj_k,proj_v,proj_o,
up_proj,down_proj,gate_proj,
lm_head
@@ A.3.3 Rank Factor Robustness
Sensitivity-based
Localization (LoSiA)
$41.70$ $\mathbf{52.23}$ $\mathbf{50.89}$ $\mathbf{36.82}$ $\mathbf{44.95}$
Gradient-based
Localization (GL)
$\mathbf{42.64}$ $51.41$ $50.62$ $36.22$ $44.88$
@@ A.3.4 Continue Learning
To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks by the order HellaSwag, PIQA, BoolQ, SIQA and Winogrande. Learning rate for LoRA is 1e41e-4 and for LoSiA is 5e55e-5. The remaining hyperparameters are consistent with Table 9. LoRA modules are merged into the backbone before subsequent task adaptation.
@@
[…]
@@ A.4.1 Memory Estimate
#Total $8LKrdb$
$2(LKR^{2}b+Vdb)$
$\max\{d^{2}b,Vdb\}$
+2LKRdb+2LKRdb
$2(LKd^{2}p^{2}b+Vdp_{o}b)$
$\max\{d^{2}b,Vdb\}$
$2Kd^{2}b$
@@ A.4.2 Latency Measurement
We measure the training latency ($\mu s$ / token) fine-tuning with different PEFT methods on LLaMA-2 7B, and the results are shown in Table 15. The experiments are conducted with $\text{cutoff\_len}=2048$ and $\text{batch \_size}=4$.